a) Find the electric potential of a point located at a distance of 0.8 m from a charge of q = -8mC.

b) Find the magnitude of the force exerted by the field E over a particle with a charge of q=+9x10-9C




1
Expert's answer
2021-10-28T15:21:35-0400

a)

"\varphi=-\frac{kq}r=90~MV,"

b)

"F=\frac{k|q_1||q_2|}{r^2}=1.01~N."